OpenStudy (anonymous):

If 2x − 2y = 10 and x + y = 5 then x equals pls show steps If 2x − 2y = 10 and x + y = 5 then x equals pls show steps @Mathematics

OpenStudy (anonymous):

2x - 2y = 10 2x = 10 + 2y x = 2y + 10 / 2 = y + 5 Replace x in either of the equations and you will get the answer.

OpenStudy (anonymous):

hi da pratu

OpenStudy (anonymous):

hullo

OpenStudy (anonymous):

prepared for maths?

OpenStudy (anonymous):

would u mind giving me a few medals?

OpenStudy (anonymous):

pratu so is the final answer 5?

OpenStudy (anonymous):

I'm not getting the answer if I replace the value of x, I think I made a mistake.....

OpenStudy (anonymous):

No, you need to find the value of x AND y....

OpenStudy (anonymous):

there must be some other way....

jhonyy9 (jhonyy9):

x=5-y

OpenStudy (anonymous):

johnny what are your steps

OpenStudy (anonymous):

we should keep x=5-y and substitute and do

OpenStudy (anonymous):

But I'm getting -4y = 0.

OpenStudy (anonymous):

Do you know the answer jhonny

OpenStudy (anonymous):

even me

OpenStudy (anonymous):

interesting question....

OpenStudy (anonymous):

I've solved this type of questions before but this seams to be tricky.

OpenStudy (anonymous):

2(5-y)-2y=10 10-2y-2y=10 -4y=0 y=0 x=y+5 x=5

OpenStudy (anonymous):

oh I see...

OpenStudy (anonymous):

I think its correct

OpenStudy (anonymous):

bulbs pratu

OpenStudy (anonymous):

no offense

OpenStudy (anonymous):

we have to find only x right

OpenStudy (anonymous):

yes da

OpenStudy (anonymous):

Adithya is correct for the first time in his life.....what an achievement!!!

OpenStudy (anonymous):

please, is easy. X=5.